# On an entropy inequality for quadratic forms and applications

### v1

@@ 3 Discretized entropy Balog-Szemerédi-Gowers theorem
Second, we take the functions $f,g$ in Theorem 3.1 to be $f(x)=\alpha x^{2}$ and $g(y)=\beta y^{2}$. By the construction $(X_{1},Y_{1})\equiv(X_{0},Y_{0}||X_{0}|>\eta,|Y_{0}|>\eta)$, the ranges of $X^{\prime\prime}$ and $Y^{\prime\prime}$ are contained in $[-2c_{1},-\eta]\cup[\eta,2c_{1}]$ with probability $1$. Now, by Example 2.13, the random variables $f(X^{\prime\prime})$ and $g(Y^{\prime\prime})$ are weakly $O_{\phi}(\max\{\eta^{-1},c_{1}\})$-bi-Lipschitz functions of $X^{\prime\prime}$ and $Y^{\prime\prime}$, respectively. It follows that
$H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}\leq 7\log K+O_{\phi,c_{1},\eta}(1).$ (44)
@@
$\begin{split}\big{(}X^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(Y^{\prime})=y)\right)\right)$-Frostman},\\ \big{(}Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s,O\left(C^{2}\cdot 2^{-ns}/P_{E}(D_{n}(X^{\prime})=x)\right)\right)$-Frostman},\end{split}$ (45)
@@
$\qquad=\frac{P_{E}(X_{1}\in B(z,r),D_{n}(Y_{1})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}\leq\frac{2P(X-pY\in B(z,r),D_{n}(X+Y)=y)}{P_{E}(D_{n}(Y^{\prime})=y)}.$
Denote by $\mu_{X}$, $\mu_{Y}$ the measures associated to the independent random variables $X$ and $Y$, respectively. Then $\mu_{X}$ and $\mu_{Y}$ are $(s,C)$-Frostman. Furthermore, define $I_{n}(y):=[y,y+2^{-n})$. We have that
[…]
The value inside the integral can be positive only when $I_{n}(y-\tilde{y})\cap B(z+p\tilde{y},r)\not=\emptyset$. Or equivalently, $I_{n}(y)\cap B(z+(p+1)\tilde{y},r)\not=\emptyset$. That is to say,
$z+(p+1)\tilde{y}-r\leq y+2^{-n},\quad z+(p+1)\tilde{y}+r\geq y.$
We can locate such $\tilde{y}$ in an interval $J$ of length at most $2r+2^{n}$. We now estimate the above expression in different ways depending on how $r$ relates to $2^{-n}$. First, for $r\geq 2^{-n}$, and recalling that $\mu_{X}$ and $\mu_{Y}$ are $(s,C)$-Frostman, we have
[…]
Next, for $r<2^{-n}$, we similarly find
[…]
This completes the proof for the claims given in (45)
As the expressions to follow get rather lengthy, we now introduce some shorthand. In what follows, $H_{n}(~{}\cdot~{}|x,y)$ is an abbreviation of $H(D_{n}(\cdot)~{}|D_{n}(X^{\prime},Y^{\prime})=(x,y)).$ By Lemma B.2, there exist random variables, depending on $(x,y)$, that have the same distributions with $(X^{\prime\prime},Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y))$, respectively. These random variables take the same value as $(X^{\prime\prime},Y^{\prime\prime})$ does. By construction, their ranges are contained in $\big{(}[-2c_{1},-\eta]\cup[\eta,2c_{1}]\big{)}^{2}$. Also note that $(X^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ and $(Y^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ are independent. Recall that the constant $\epsilon_{1}=\epsilon_{1}(s)$ in Lemma 4.1 is independent of the concrete random variables. Finally, recall (45). Now we can apply Lemma 4.1 with these random variables, with $C_{1}=O_{C}\big{(}2^{-ns}/P_{E}(D_{n}(Y^{\prime})=y)\big{)}$, $C_{2}=O_{C}\big{(}2^{-ns}/P_{E}(D_{n}(X^{\prime})=x)\big{)},$ and $c_{2}=\eta$, to deduce that
@@
$\qquad+O\big{(}-2ns-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}+O_{\phi,s,C,c_{1},\eta}(1).$ (46)
By conditional independence, and combining Corollary 2.22, we also have that
@@
Using a similar argument, Lemma B.2, and Example 2.17, one obtains
@@
$=\max\{H_{n}(X^{\prime\prime}|x,y),H_{n}(Y^{\prime\prime}|x,y)\}+O_{\phi,c_{1},\eta}(1).$
@@
$\begin{split}&\min\{H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y),H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\}\geq\frac{H_{n}(X^{\prime\prime}|x,y)+H_{n}(Y^{\prime\prime}|x,y)}{2}+O_{\phi,c_{1},\eta}(1).\end{split}$ (47)
@@
$\quad\geq n(s+\epsilon_{1})+O\big{(}-2ns-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}+O_{\phi,s,C,c_{1},\eta}(1).$
@@
$\qquad\qquad\geq n(s+\epsilon_{1})+O(-2ns+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))+O_{\phi,s,C,c_{1},\eta}(1).$ (48)
@@
$\begin{split}&\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)+\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)\\ &\qquad-\frac{(H_{n}(X_{1})-\log K)+(H_{n}(Y_{1})-\log K)}{2}-ns+O(H_{n}(X_{1})+H_{n}(Y_{1})-2ns)\\ &\qquad\geq n\epsilon_{1}+O_{\phi,s,C,c_{1},\eta}(1).\end{split}$ (49)
Recalling (40) and the notation $\log K=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1)$, the left-hand side of (49) is
$\leq O\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}-ns\right)+15\log K=O(\Psi)+O_{C}(\sqrt{\eta}n)+O_{C}(1).$
@@
$\tilde{C}\Psi+O_{C}(\sqrt{\eta}n)\geq n\epsilon_{1}+O_{\phi,s,C,c_{1},\eta}(1).$
Equivalently,
$\Psi\gg n\epsilon_{1}+O_{C}(\sqrt{\eta}n)+O_{\phi,s,C,c_{1},\eta}(1).$
[…]
Now we are ready to finish the proof of Theorem 1.1.
Proof of Theorem 1.1.
For $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$, the non-degeneracy gives that $a_{2}^{2}\neq 4a_{1}a_{3}$.
When $a_{3}\neq 0$, we have
$\phi(x,y)=\left(a_{1}-\frac{a_{2}^{2}}{4a_{3}}\right)x^{2}+a_{3}\left(\frac{a_{2}}{2a_{3}}x+y\right)^{2}.$
Take $\tilde{X}=\frac{a_{2}}{2a_{3}}X$ and $\tilde{Y}=Y$. Then $\phi(X,Y)=\tilde{\phi}(\tilde{X},\tilde{Y})$, where
$\tilde{\phi}(\tilde{x},\tilde{y})=\left(\frac{4a_{1}a_{3}^{2}}{a_{2}^{2}}-a_{3}\right)\tilde{x}^{2}+a_{3}(\tilde{x}+\tilde{y})^{2}.$
When $a_{1}\neq 0$, similar arguments lead to $\phi(X,Y)=\tilde{\phi}(\tilde{X},\tilde{Y})$, with $\tilde{X}=\frac{a_{2}}{2a_{1}}Y$, $\tilde{Y}=X$, and
$\tilde{\phi}(\tilde{x},\tilde{y})=\left(\frac{4a_{1}^{2}a_{3}}{a_{2}^{2}}-a_{1}\right)\tilde{x}^{2}+a_{1}(\tilde{x}+\tilde{y})^{2}.$
When $a_{1}=a_{3}=0$, one has $a_{2}\neq 0$ and $\phi(X,Y)=\tilde{\phi}(\tilde{X},\tilde{Y})$, where $\tilde{X}=X$, $\tilde{Y}=Y$, and
$\tilde{\phi}(\tilde{x},\tilde{y})=-\frac{a_{2}}{4}(\tilde{x}-\tilde{y})^{2}+\frac{a_{2}}{4}(\tilde{x}+\tilde{y})^{2}.$
In all cases, the random variables $\tilde{X}$ and $\tilde{Y}$ are independent, and both are $(s,O_{\phi,C}(1))$-Frostman. The ranges of $\tilde{X}$ and $\tilde{Y}$ are both contained in $[-\widetilde{c_{1}},\widetilde{c_{1}}]$, with $\widetilde{c_{1}}=O_{\phi,c_{1}}(1)$.
Applying Lemma 5.1 with the quadratic form $\tilde{\phi}(\tilde{x},\tilde{y})$ and the random variables $\tilde{X}$ and $\tilde{Y}$, we obtain that
$\max\{H_{n}(\xi X+\zeta Y),\,H_{n}(\phi(X,Y))\}\geq n(s+\varepsilon_{2})$
for some non-zero rational numbers $\xi=\xi(\phi)$ and $\zeta=\zeta(\phi)$. Let $\xi=\frac{k}{m}$ and $\zeta=\frac{l}{m}$, with $m$ being the greatest common divisor of $\xi$ and $\zeta$. By Lemmas 2.23 and 2.26, there are some $d_{1}=d_{1}(\xi,\zeta)$ and $d_{2}=d_{2}(\xi,\zeta)$ such that
$H_{n}(\xi X+\zeta Y)$ $\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-d_{1}H_{n}(X)-d_{2}H_{n}(Y)+O_{\xi,\zeta}(1)$
$\leq(d_{1}+d_{2}+1)H_{n}(X+Y)-(d_{1}+d_{2})(ns-\log C)+O_{\phi}(1).$
That is to say,
$H_{n}(X+Y)\geq\frac{(d_{1}+d_{2})ns+H_{n}(\xi X+\zeta Y)}{d_{1}+d_{2}+1}+O_{\phi,C}(1).$
Now the theorem follows by taking $\epsilon=\frac{\epsilon_{2}}{2(d_{1}+d_{2}+1)}$ when $n$ is sufficiently large.

### v2

@@ 4.2 Diagonal forms with dependent random variables
Second, we take the functions $f,g$ in Theorem 3.1 to be $f(x)=\alpha x^{2}$ and $g(y)=\beta y^{2}$. By the construction $(X_{1},Y_{1})\equiv(X,Y||(X,Y)|>\eta)$, the ranges of $X^{\prime\prime}$ and $Y^{\prime\prime}$ are contained in $c,-\eta/\sqrt{2}]\cup[\eta/\sqrt{2},c]$ with probability $1$. Now, by Example 2.12, the random variables $f(X^{\prime\prime})$ and $g(Y^{\prime\prime})$ are weakly $O_{\phi}(\max\{\eta^{-1},c\})$-bi-Lipschitz functions of $X^{\prime\prime}$ and $Y^{\prime\prime}$, respectively. It follows that
$H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|X^{\prime},Y^{\prime})-\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}\leq 7\log K+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$ (46)
@@
$\begin{split}\big{(}X^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s_{1},O\left(C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\right)\right)$-Frostman},\\ \big{(}Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y)\big{)}\text{ is $\left(s_{2},O\left(C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\right)\right)$-Frostman},\end{split}$ (47)
@@
$\qquad=\frac{P_{E}(X_{1}\in B(z,r),D_{n}(Y_{1})=y)}{P_{E}(D_{n}(Y^{\prime})=y)}\leq\frac{2P(X\in B(z,r),D_{n}(Y)=y)}{P_{E}(D_{n}(Y^{\prime})=y)}.$
Since $(X,Y)$ is jointly $(s_{1},s_{2},C)$-Frostman, we have
$P(X\in B(z,r),D_{n}(Y)=y)\ll Cr^{s_{1}}2^{-ns_{2}}.$
This completes the proof for the claims given in (47). We mention that these two Frostman constants in (47) are $\gg 1$.
As the expressions to follow get rather lengthy, we now introduce some shorthand. In what follows, $H_{n}(~\cdot~|x,y)$ is an abbreviation of $H(D_{n}(\cdot)~|D_{n}(X^{\prime},Y^{\prime})=(x,y)).$ By Lemma A.2, there exist random variables, depending on $(x,y)$, that have the same distributions with $(X^{\prime\prime},Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y))$, respectively. These random variables take the same value as $(X^{\prime\prime},Y^{\prime\prime})$ does. By construction, their ranges are contained in $\big{(}[-c,-\eta/\sqrt{2}]\cup[\eta/\sqrt{2},c]\big{)}^{2}$ with probability $1$. Also note that $(X^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ and $(Y^{\prime\prime}|D_{n}(X_{1},Y_{1})=(x,y))$ are independent. Recall that the constant $\epsilon_{1}=\epsilon_{1}(s)$ in Lemma 4.1 is independent of the concrete random variables. Finally, recall (47). Now we can apply Lemma 4.1 with these random variables, with $C_{1}=O\big{(}C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\big{)}$, $C_{2}=O\big{(}C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\big{)}$, $c_{1}=c$ and $c_{2}=\eta$, to deduce that
@@
$\qquad+O\big{(}\log_{+}C-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}$
$\qquad+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1)$ (48)
for some $\epsilon_{1}=\epsilon_{1}(s)>0$. By conditional independence, and combining (13), we also have that
@@
Using a similar argument, Lemma A.2, and Example 2.16, one obtains
@@
$\qquad=\max\{H_{n}(X^{\prime\prime}|x,y),H_{n}(Y^{\prime\prime}|x,y)\}+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$
@@
$\min\{H_{n}(X^{\prime\prime}+Y^{\prime\prime}|x,y),H_{n}(\alpha X^{\prime\prime 2}+\beta Y^{\prime\prime 2}|x,y)\}$
$\qquad\geq\frac{H_{n}(X^{\prime\prime}|x,y)+H_{n}(Y^{\prime\prime}|x,y)}{2}+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi}(1).$ (49)
@@
$\quad\geq n(s+\epsilon_{1})+O\big{(}-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big{)}$
$\qquad\qquad\qquad\quad+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
@@
$\qquad\qquad\geq n(s+\epsilon_{1})+O(-n(s_{1}+s_{2})+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))$
$\qquad\qquad\qquad\quad+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$ (50)
@@
$\begin{split}&\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)+\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}+7\log K\right)\\ &\qquad-\frac{(H_{n}(X_{1})-\log K)+(H_{n}(Y_{1})-\log K)}{2}-ns+O(H_{n}(X_{1})+H_{n}(Y_{1})-n(s_{1}+s_{2}))\\ &\qquad\geq n\epsilon_{1}+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).\end{split}$ (51)
Recalling (43), the fact $s_{1}+s_{2}\geq 2s$, and the notation $\log K=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(1)$, the left-hand side of (51) is
$\leq O\left(\frac{H_{n}(X_{1})+H_{n}(Y_{1})}{2}-ns\right)+15\log K$
$=O(\Psi)+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
@@
$\tilde{C}\Psi\geq n\epsilon_{1}+O(C\eta^{2s}n)+O(\log_{+}C)+O(\log_{+}c)+O(\log(1/\eta))+O_{\phi,s}(1).$
Recall that we require $0<\eta<\min\{2^{-1}(1/2C)^{1/2s},1/10\}$, $0<\epsilon_{1}<1$, and $\epsilon_{1}$ only depends on $s$. Now assume that the implied constant in the term $O(C\eta^{2s}n)$ is $C^{\prime}$, with $C^{\prime}\geq 1$. We then take $\eta=\min\Big{\{}\frac{1}{2}\big{(}\frac{\epsilon_{1}}{2C^{\prime}C}\big{)}^{1/2s},\,\frac{1}{10}\Big{\}}$. Then $C^{\prime}C\eta^{2s}n\leq n\epsilon_{1}/2$. Therefore,
$\Psi\gg n\epsilon_{1}+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1).$
The proof is completed. ∎
Next, we prove Theorem 1.3 in the following form.
Theorem 1.3’. Let $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $0<s<1$ and $C_{1},C_{2},c>0$. There is a positive constants $\epsilon_{3}=\epsilon_{3}(s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
(i) For all $n\in\mathbb{N}$,
$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\geq n(s+\epsilon_{3})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1).$
(ii) For all $n\in\mathbb{N}$,
$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(XY)\right\}\geq n(s+\epsilon_{3})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1).$
Proof.
(i) Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.22 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman. Applying Theorem 1.4’, there is some $\epsilon_{2}=\epsilon_{2}(s)$ such that
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}$
$\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1)$
for all $n\in\mathbb{N}$. In view of $s_{1}+s_{2}\geq 2s$, Lemma 2.24(ii) shows that
$H_{n}(X,Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+ns+O(\log_{+}C_{1}C_{2})+O(1).$
The conclusion then follows with $\epsilon_{3}=\epsilon_{2}/2$.
(ii) Notice that
$xy=\frac{1}{4}(x+y)^{2}-\frac{1}{4}(x-y)^{2}.$
Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.23 it is also jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O(1)$.
Applying Theorem 1.4’ with $\alpha=1/4$, $\beta=-1/4$ and the random variable (X+Y,XY)(X+Y,X-Y), we obtain
$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(2X),\,H_{n}(XY)\right\}$ (52)
$\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$
for some $\epsilon_{2}=\epsilon_{2}(s)$. If we switch the role of $X$ and $Y$, then similar conclusion holds, with $H_{n}(2X)$ replaced by $H_{n}(2Y)$. Therefore, the left-hand side of (52) can be replaced by
$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,\frac{H_{n}(2X)+H_{n}(2Y)}{2},\,H_{n}(XY)\right\}.$
By Lemma 2.24, combining $s_{1}+s_{2}\geq 2s$, the second item satisfies
$H_{n}(X,Y)\geq\frac{H_{n}(2X)+H_{n}(2Y)+O(1)}{2}+ns+O(\log_{+}C_{1}C_{2})+O_{s}(1).$
Combining the fact
$H_{n}(X,Y)=H_{n}(X+Y,X-Y)+O(1)\leq H_{n}(X+Y)+H_{n}(X-Y)+O(1),$
the first conclusion follows by taking $\epsilon_{3}=\epsilon_{2}/2$. ∎
4.3 The general forms
Now we are ready to prove Theorem 1.2. It actually holds for any non-degenerate real quadratic form $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ that is congruent over $\mathbb{Q}$ to a non-degenerate real diagonal form, i.e.,
$\phi(x,y)=\alpha(\xi_{11}x+\xi_{12}y)^{2}+\beta(\xi_{21}x+\xi_{22}y)^{2}$ (53)
for some $\alpha,\beta\in{\mathbb{R}}^{\ast}$ and $\xi_{11},\xi_{12},\xi_{21},\xi_{22}\in\mathbb{Q}$ with $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$. Here, all the constants $\alpha,\beta,\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ depends on the coefficients $a_{1},a_{2},a_{3}$ in $\phi(x,y)$.
Theorem 1.2’ Let $\phi(x,y)$ be in the form (53). Let $0<s<1$, $C\geq 1$ and $c>0$. There is a positive constant $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables with ranges contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(x,y))\right\}\geq n(s+\epsilon)+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1)$
for all $n\in\mathbb{N}$.
Proof.
By Lemma 2.23, the random variable $Z:=(\xi_{11}X+\xi_{12}Y,\xi_{21}X+\xi_{22}Y)$ is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, where $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$. Moreover, the range of $Z$ is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$ with $c^{\prime\prime}=O_{\phi}(c)$. Applying Theorem 1.4’ with $Z$, we obtain some positive number $\epsilon_{2}=\epsilon_{2}(s)$ such that
$\max\left\{\frac{H_{n}(\xi_{11}X+\xi_{12}Y)+H_{n}(\xi_{21}X+\xi_{22}Y)}{2},\,H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)},\,H_{n}\big{(}\phi(X,Y)\big{)}\right\}$
$\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$ (54)
for all $n\in\mathbb{N}$.
Next, using Lemma 2.21, let us assume that $d=d(\xi_{11},\xi_{12},\xi_{21},\xi_{22})=d(\phi)$ is the positive number satisfying
$H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}\leq dH_{n}(X+Y)-(d-1)\frac{H_{n}(X)+H_{n}(Y)}{2}+O(1).$ (55)
Combining Lemma 2.24(i) and the fact that $s_{1}+s_{2}\geq 2s$, we see that the right-hand side of (55) is
$\leq dH_{n}(X+Y)-(d-1)\frac{(ns_{1}-\log C_{1}+O(1))+(ns_{2}-\log C_{2}+O(1))}{2}+O(1)$
$\leq dH_{n}(X+Y)-(d-1)ns+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1).$
This leads to
$H_{n}(X+Y)\geq\frac{H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}+(d-1)ns}{d}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1).$
Similarly, we can deduce that
$H_{n}(X+Y)\geq\frac{H_{n}(\xi_{11}X+\xi_{12}Y)+(d^{\prime}-1)ns}{d^{\prime}}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1),$
and also
$H_{n}(X+Y)\geq\frac{H_{n}(\xi_{21}X+\xi_{22}Y)+(d^{\prime\prime}-1)ns}{d^{\prime\prime}}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1)$
for some positive numbers $d^{\prime}=d^{\prime}(\phi)$ and $d^{\prime\prime}=d^{\prime\prime}(\phi)$. Now we deduce from (54) that
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{5})+O_{\phi,s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$
for $\epsilon_{4}=\epsilon_{2}/(d+d^{\prime}+d^{\prime\prime})$ and all $n\in\mathbb{N}$. Theorem 1.2 then follows. ∎